OpenStudy (anonymous):

What is the solution set of 4x = y and 2x2 − y = 0? {(0.5, -0.5), (2, -8)} {(0, 0), (2, 8)} {(1.5, 2), (8, 2)} {(-1.5, -2), (2, 8)}

OpenStudy (anonymous):

substitute for y in the second equation, since it's already isolated in the first eq'n. 2x^2 - (4x) = 0 2x (x - 2) = 0 set the "brackets" = 0. 2x = 0, therefore x=0 x-2=0, therfore x=2

OpenStudy (anonymous):

the 2x2 (the last 2 is squared)

OpenStudy (anonymous):

\[2*2^{2} \]

OpenStudy (anonymous):

Yes like that

OpenStudy (anonymous):

I dont get what you said @Nerd94

OpenStudy (anonymous):

X=0 and X=2

OpenStudy (anonymous):

x=0 y=2

OpenStudy (anonymous):

Oh so its B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thats right! Thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

gotta retake it tho so i still need help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

An archway is modeled by the equation y = -2x2 + 8x. A rod is to be placed across the archway at an angle defined by the equation x − 2.23y + 10.34 = 0. If the rod is attached to the archway at points A and B, such that point B is at a higher level than point A, at what distance from the ground level is point B?

OpenStudy (anonymous):

all the ???????? is meant to be a minus/subtraction sign (-)

OpenStudy (anonymous):

8 units 6 units 5 units 3 units

OpenStudy (anonymous):

@Junie.Michell98

OpenStudy (anonymous):

@Nerd94

OpenStudy (anonymous):

if Im not mistaken i should be 8 units

OpenStudy (anonymous):

8 units

OpenStudy (anonymous):

It was wrong

OpenStudy (anonymous):

oh im sorry

OpenStudy (anonymous):

gotta get the next 4 right or i fail again

OpenStudy (anonymous):

A fountain on a lake sprays water in a parabolic arch modeled by the equation y = −0.4x2 + 3x, where y is the height in feet and x is the horizontal distance. A beam of light is passed through the fountain to create a rainbow effect. If the beam is directed at an angle modeled by the equation -2.2x + 4.9y = 8.18, at what distance from the ground will the beam first touch the water spray?

OpenStudy (anonymous):

the answer choices... 0.6 feet 2 feet 4.2 feet 5.6 feet

OpenStudy (anonymous):

@Demonx341

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

help with question above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@iGreen will u explain how to do this to @SavannahPaige18 ??

OpenStudy (anonymous):

sorry but i dont know how to do this that is why i asked igreen.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@Nerd94

OpenStudy (anonymous):

4.2 ft.

OpenStudy (anonymous):

It was wrong

OpenStudy (anonymous):

What is the solution set of y = x2 + 2x + 7 and y = x + 7? {(0, 7), (-1, 6)} {(0, 7), (-7, 0)} {(0, 7), (1, 8)} {(-2, 0), (4, 0)} ITS NOT B****